OpenStudy (anonymous):

John invests $10,000 for two years at 10% compounded annually. How much will John have after the two years?@britbrat4290

OpenStudy (anonymous):

A = 10,000 (1 + 0.1) ^ 2 $12,100

hero (hero):

The formula for annual compounding is: \[A = P(1 + r)^t\] A=Account Balance P=Principal r=rate (decimal form) t=time (in years)

OpenStudy (anonymous):

thanks @britbrat4290

OpenStudy (anonymous):

no prob :)